OpenStudy (anonymous):

find the value of θ in the equation cos2θ+sin²θ=cos³θ+3θ-6

OpenStudy (anonymous):

Well cos(2t)=cos^2(t)-sin^2(t); so the sin^2(t)'s cancel.

OpenStudy (anonymous):

okay so i did that and got cos^2x= cos^3+3x-6 then cos^2x-cos^3x = 3x-6

OpenStudy (anonymous):

@malevolence19